Gaucho Services starts life with all-equity financing and a cost of equity of 14%. Suppose it refinances to the following market-value capital structure:

Debt (D) 45% at rD = 9.4%
Equity (E) 55%


Use MM’s proposition 2 to calculate the new cost of equity. Gaucho pays taxes at a marginal rate of Tc = 30%. Calculate Gaucho’s after-tax weighted-average cost of capital. (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)Gaucho Services starts life with all-equity financing and a cost of equity of 14%. Suppose it refinances to the following mar
Answer #1
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 14+0.818181818181818*(14-9.4)*(1-0.3)
Levered cost of equity = 16.63
Weight of equity = 1-D/A
Weight of equity = 1-0.45
W(E)=0.55
Weight of debt = D/A
Weight of debt = 0.45
W(D)=0.45
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 9.4*(1-0.3)
= 6.58
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=6.58*0.45+16.63*0.55
WACC =12.11%